# Chiara Arina Centre for Cosmology, Particle Physics and Phenomenology (CP3), Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

### v1

@@ III.3.4 Gaps in coverage
[…]
@@
The quark-philic scenario discussed above provides an example of this gap in coverage and how it could be addressed with minor modifications to existing search strategies. For instance, we could consider the ATLAS displaced vertex search [212] (ATLAS-SUSY-2016-08), which relies on a $E_{T}^{\mathrm{miss}}$ trigger and is sensitive to decay lengths between 4 mm and 30 cm. The signal region defined in this search requires at least one displaced vertex with five or more tracks and an invariant mass exceeding 10 GeV. In the left panel of figure 35, we present the distribution of the number of displaced vertices as a function of the number of tracks $n_{\rm trk}$ associated with it and of the invariant mass oof the vertex $m_{\rm DV}$, in the framework of a quark-philic CDFO benchmark model with $M_{X}=450$ GeV, $\Delta m=31$ GeV, and $c\tau=0.85$ cm. Most vertices have $93<n_{\rm trk}<$ 9 and $m_{\rm DV}<6$ GeV, making them too soft to satisfy the ATLAS analysis requirements. For this benchmark, the $m_{\rm DV}$ cut reduces the signal efficiency from approximately $6\%$ to $0.1\%$, therefore significantly diminishing the search sensitivity. Although the high $m_{\rm DV}$ threshold effectively suppresses SM backgrounds, it could likely be lowered without a substantial increase in background contamination. Indeed, the ATLAS search observed zero vertices with $n_{\rm trk}>5$ and $m_{\rm DV}>4$ GeV so that lowering the $m_{\rm DV}$ requirement, for instance to 5 GeV, could still maintain an effective SM background suppression. Assuming that the SM background associated with the modified selection $m_{\rm DV}>5$ GeV remains negligible (i.e. fewer than one vertex), we estimate in the right panel of figure 35 the sensitivity improvement from reducing the invariant mass threshold from 10 GeV (red exclusion curve) to 5 GeV (orange exclusion curves), as further detailed in [259]. This demonstrates that a significant portion of parameter space becomes accessible when mildly soft displaced vertices are included, this region being in addition not covered by other LLP searches.
@@ IV.2 Dark matter relic density
In order to write (35), we have assumed that all particles involved in (co-)annihilation and conversion processes are in kinetic equilibrium at the time of DM production. For non-thermal DM production discussed in sections IV.2.3 and IV.2.4, the relevant relation (42) can be obtained from (35) without making any assumptions on the DM distribution, but after instead enforcing a negligible initial abundance for $X$ and suppressed couplings to $Y$. Equation (35) is thus also perfectly suitable for evaluating the DM comoving number density $Y_{X}$ in FI and SW scenarios for which the loss terms for the dark matter particle $X$ are neglected. [^12] For further discussion on kinetic decoupling, see e.g. [279, 280, 34]. In addition, the semi-classical Boltzmann equation (35) neglects various quantum and thermal effects, such as quantum coherence, quantum statistics, screening, multiple scatterings, and particle interactions with coherent condensates. A systematic framework for deriving quantum kinetic equations that incorporate these effects in relic density computations is provided by the Schwinger-Keldysh formalism (see, e.g., [281, 282, 283, 284, 285, 286, 287], as well as [288, 289, 290, 291] for related computations in equilibrium and [286] for a lattice-based approach). In particular for freeze-in scenarios quantum and thermal effects can alter the relic density significantly. This was demonstrated in [289] for a Majorana DM model and in [284] for a scalar DM model, respectively. The latter work, which employs non-equilibrium quantum field theory from first principles, quantifies the expected corrections to various approaches commonly used in the literature (freeze-in with only decays, including scattering with and without thermal masses, hard thermal loop approximation, etc.) and provides recommendations for accurate results in phenomenological studies. For freeze-out scenarios and in particular for co-annihilating configurations, significant impact is expected from the Sommerfeld effect and bound state formation, as discussed in section IV.2.5. However, only small corrections are expected due to thermal effects, as detailed in the context of dark matter annihilation in, e.g., [292, 281] and Sommerfeld and bound state formation in, e.g., [288, 293, 294, 295, 296, 297].
@@
$\begin{split}\gamma_{ij\to kl}=&\int\mathrm{d}\phi_{i}\,\mathrm{d}\phi_{j}\,f_{i}^{\rm eq}(t,p_{i})\,f_{j}^{\rm eq}(t,p_{j})\int\mathrm{d}\phi_{k}\,\mathrm{d}\phi_{l}\,(2\pi)^{4}\delta^{4}(p_{i}\!+\!p_{j}\!-\!p_{k}\!-\!p_{l})\,\big{|}{\cal M}_{ij\to kl}\big{|}^{2}\,,\\ \gamma_{i\to jk}=&\int\mathrm{d}\phi_{i}\,\mathrm{d}\phi_{j}\,\mathrm{d}\phi_{k}\,f_{i}^{\rm eq}(t,p_{i})\,(2\pi)^{4}\delta^{4}(p_{i}-p_{j}-p_{k})\,\big{|}{\cal M}_{i\to jk}\big{|}^{2}\,,\end{split}$ (36)
@@ IV.2.1 Canonical freeze-out
$\begin{split}g_{\star,\text{eff}}^{\frac{1}{2}}=&\ \frac{g_{\star S}}{\sqrt{g_{\star}}}\left(1+\frac{T}{3g_{\star S}}\frac{\mathrm{d}g_{\star S}}{\mathrm{d}T}\right)\,,\\ \left\langle\sigma_{\text{eff}}v\right\rangle=&\ \sum_{i,j}\left\langle\sigma_{ij}v\right\rangle\frac{Y_{i}^{\rm eq}Y_{j}^{\rm eq}}{\tilde{Y}^{\rm eq}\tilde{Y}^{\rm eq}}\,,\end{split}$ (40)
@@
Figure 38: Representative LO Feynman diagrams contributing to the DM annihilation cross section $\langle\sigma_{\text{eff}}v\rangle$. They include $t$-channel DM annihilations (top left), $XY$ co-annihilations (top central and right), and mediator pair-annihilations (bottom row) with the blob accounting for channels yielding SM electroweak bosons or gluons. The final state $q$ represents any quark flavour, and may include a sum over flavours depending on the model. For leptophilic models, $q$ can be replaced by a lepton $\ell$, and for models with self-conjugate DM, additional $YY$ annihilation diagrams must be added. Moreover, in cases where the S-wave contribution to $\langle\sigma_{\text{ann}}v\rangle(X\bar{X}\to q\bar{q})$ is helicity suppressed, NLO corrections become relevant (see the diagrams in figure 40).
@@ IV.2.2 Conversion-driven freeze-out
$\begin{split}\frac{\mathrm{d}Y_{X}}{\mathrm{d}x}=&\ \frac{1}{Hxs}\left[\gamma_{Y\to X}\left(\frac{Y_{Y}}{Y_{Y}^{\rm eq}}-\frac{Y_{X}}{Y_{X}^{\rm eq}}\right)+\gamma_{\bar{Y}\to X}\left(\frac{Y_{\bar{Y}}}{Y_{\bar{Y}}^{\rm eq}}-\frac{Y_{X}}{Y_{X}^{\rm eq}}\right)\right]\,,\\ \frac{\mathrm{d}Y_{Y}}{\mathrm{d}x}=&\ \frac{-1}{Hxs}\left[\gamma_{Y\bar{Y}}\left(\frac{Y_{Y}Y_{\bar{Y}}}{Y_{Y}^{\rm eq}Y_{\bar{Y}}^{\rm eq}}-1\right)+\gamma_{Y\to X}\left(\frac{Y_{Y}}{Y_{Y}^{\rm eq}}-\frac{Y_{X}}{Y_{X}^{\rm eq}}\right)\right]\,,\end{split}$ (41)
@@ IV.2.3 Freeze-in
Thermal corrections, such as introducing thermal masses for vector bosons such as gluons or the electroweak bosons, should be included in the DM abundance calculation. They imply in particular that decay processes typically dominate freeze-in production [252, 204]. Scattering processes such as $VY\to Xf$ may contribute slightly to the relic abundance, up to $\sim 25\%$, when $V$ is a gluon and $f$ is a coloured SM fermion. A more precise treatment of such thermal effects, particularly for small mass splittings, may nevertheless be necessary [289]. For simplicity, we focus here on mediator decay contributions during a radiation-dominated era to illustrate the main aspects of freezing-in DM production. In this case, (42) reduces to
@@ IV.2.5 Non-perturbative effects
The freeze-out of dark sector annihilations typically occurs when the annihilating particles are non-relativistic. This regime allows for significant non-perturbative effects, namely the Sommerfeld Effect (SE) and Bound State Formation (BSF), that affect the annihilation cross section through multiple exchanges of light states between the initial or final state particles. Non-perturbative corrections in the final state are neglected here, since the annihilation products are much lighter than the dark sector particles so that they are relativistic.
Figure 39: Diagramatic representation of bound state ($\cal B$) formation of two coloured particles via the emission of a gluon, and of the Sommerfeld Effect relevant for the initial state. Figure adapted from [301].
Simplified $t$-channel DM models always involve a dark sector particle $Y$ carrying SM charges. Consequently, annihilation involving two SM-charged dark sector particles is subject to corrections from multiple exchanges of (massless) gauge bosons. Each exchange of a massless gauge boson contributes parametrically with a factor of $\alpha/v$ where $\alpha$ is the relevant coupling constant and $v$ is the relative velocity between the two initial-state particles, which could be sizeable if $\alpha\sim v$. During the freeze-out of DM annihilations, $v\sim 0.1$. This thus necessitates the resummation of the ladder diagrams shown in figure 39 if $\alpha\sim 0.1$, which is particularly relevant for mediator particles $Y$ charged under $SU(3)_{C}$ (as in the models considered). In the non-relativistic regime, this resummation reduces to solving the Schrödinger equation for the wave functions of the initial-state particles, incorporating a colour potential as discussed in [305, 306, 307, 308]. Since gluons are massless, the potential at energy scales far above the QCD confinement scale is Coulomb-like, and can be either attractive or repulsive depending on the representation of the initial state particles. In the context of simplified $t$-channel DM models, the dark sector particles are either singlet or lie in the (anti-)fundamental representation of $SU(3)_{C}$. Dark sector annihilation therefore leads to the following initial state colour configurations,
@@
where $\langle\sigma_{\text{BSF}}v\rangle$ is the thermally averaged cross section for bound state formation under the emission of a gluon [300]. The prefactor depends on the bound state decay and ionisation rates $\Gamma_{\text{dec}}$ and $\Gamma_{\text{ion}}$, and reflects that only bound state decays into SM particles deplete the dark sector. When excited bound states are included, transitions between different states $i$ and $j$ with rates $\Gamma_{\text{trans}}^{i\rightarrow j}$ must be considered, which yields [258, 261]
@@
When summing the bound-state formation contributions to the cross section over all possible bound states in the final state, unitarity violation however arises [312, 313].
@@
The Sommerfeld effect and bound state formation have been extensively studied in $t$-channel scenarios [288, 314, 302, 300, 303, 301, 293, 294, 258]. Their impact on the interpretation of the constraints existing on the parameter space is in particular explored in [301] for freeze-out scenarios and in [258] for CDFO models.
@@ IV.2.6 Altered cosmological histories
As shown section III.3, these findings modify the predictions for LLP signatures, demonstrating the intricate dependence of the results on the reheating potential and the main inflaton decay channel. This arises from the power-law dilution of the relic abundance,
$Y_{X}\sim\Gamma_{Y}\times\begin{cases}\left(\dfrac{T_{\text{rh}}}{M_{Y}}\right)^{4k-1}\,&\text{Bosonic Reheating,}\\ \left(\dfrac{T_{\text{rh}}}{M_{Y}}\right)^{\frac{9-k}{k-1}}\;&\text{Fermionic Reheating,}\end{cases},$ (55)
which holds if $M_{Y}\gtrsim T_{\text{RH}}$. While dilution of the relic abundance necessitates larger mediator decay widths in freeze-in scenarios, it reduces the DM-SM interaction strength for thermal freeze-out production. Since freeze-out during reheating has not been explored in the context of simplified $t$-channel DM models, we refer instead the reader to general discussions in [277].
@@ IV.3 Dark matter constraints
In this section, we describe the most relevant astrophysical and cosmological constraints that could be applied on the models studied.
In sections IV.3.1 and IV.3.2, we focus on astrophysical searches probing DM interactions at the present time, where DM is cold and clustered within halos in galaxies and galaxy clusters. The typical velocities $v$ involved range from approximately $10^{-5}c$ to $10^{-2}c$, with $v\sim 10^{-3}c$ in the Milky Way, and current detection methods primarily include indirect and direct detection approaches. Indirect detection relies on the annihilation of DM into SM particles in regions with high DM density, such as the Galactic Centre and dwarf spheroidal galaxies (dSphs). The SM particles produced through DM annihilations subsequently decay, shower, and hadronise to produce a flux of stable cosmic rays ($\gamma$, $e^{+}$, $\bar{p}$, and $\nu$), which can be detected by space-based and ground-based telescopes. Direct detection, on the other hand, relies on observing elastic scattering of DM off protons or neutrons within underground detectors, where the scattering produces a measurable recoil of the associated nucleus. It should be noted that these detection methods are most effective in probing the canonical freeze-out regime, where DM self-annihilation or DM-nucleon scattering rates are sufficiently large. Conversely, scenarios involving freeze-in, superWIMP, or conversion-driven freeze-out predict extremely small rates that are typically not testable with these methods. Instead, such scenarios often give rise to the interesting long-lived particle signals at colliders that have been discussed in section III.3.

### v2

@@ III.3.4 Gaps in coverage
The scenarios discussed in section III.3.3 illustrate the importance of LLP searches in testing various DM mechanisms. Despite the broad scope of the LLP search programme at the LHC, specific regions of the various parameter spaces considered remain uncovered by current search strategies. In particular, within the CDFO scenario, configurations with $c\tau\lesssim 1$ cm ($\Delta m\gtrsim 10$$30$ GeV) are not covered because they fall outside the lifetime range addressed by disappearing track searches, and/or the LLP visible decay products are too soft to be detected in conventional displaced vertex searches. This gap in coverage is approximately represented by the white region shown in figure 27.
@@
The quark-philic scenario discussed above provides an example of this gap in coverage and how it could be addressed with minor modifications to existing search strategies. For instance, we could consider the ATLAS displaced vertex search [212] (ATLAS-SUSY-2016-08), which relies on a $E_{T}^{\mathrm{miss}}$ trigger and is sensitive to decay lengths between 4 mm and 30 cm. The signal region defined in this search requires at least one displaced vertex with five or more tracks and an invariant mass exceeding 10 GeV. In the left panel of figure 35, we present the distribution of the number of displaced vertices as a function of the number of tracks $n_{\rm trk}$ associated with it and of the invariant mass oof the vertex $m_{\rm DV}$, in the framework of a quark-philic CDFO benchmark model with $M_{X}=450$ GeV, $\Delta m=31$ GeV, and $c\tau=0.85$ cm. Most vertices have $3<n_{\rm trk}<9$ and $m_{\rm DV}<6$ GeV, making them too soft to satisfy the ATLAS analysis requirements. For this benchmark, the $m_{\rm DV}$ cut reduces the signal efficiency from approximately $6\%$ to $0.1\%$, therefore significantly diminishing the search sensitivity. Although the high $m_{\rm DV}$ threshold effectively suppresses SM backgrounds, it could likely be lowered without a substantial increase in background contamination. Indeed, the ATLAS search observed zero vertices with $n_{\rm trk}>5$ and $m_{\rm DV}>4$ GeV so that lowering the $m_{\rm DV}$ requirement, for instance to 5 GeV, could still maintain an effective SM background suppression. Assuming that the SM background associated with the modified selection $m_{\rm DV}>5$ GeV remains negligible (i.e. fewer than one vertex), we estimate in the right panel of figure 35 the sensitivity improvement from reducing the invariant mass threshold from 10 GeV (red exclusion curve) to 5 GeV (orange exclusion curves), as further detailed in [259]. This demonstrates that a significant portion of parameter space becomes accessible when mildly soft displaced vertices are included, this region being in addition not covered by other LLP searches.
@@ IV.2 Dark matter relic density
In order to write (35), we have assumed that all particles involved in (co-)annihilation and conversion processes are in kinetic equilibrium at the time of DM production. For non-thermal DM production discussed in sections IV.2.3 and IV.2.4, the relevant relation (42) can be obtained from (35) without making any assumptions on the DM distribution, but after instead enforcing a negligible initial abundance for $X$ and suppressed couplings to $Y$. Equation (35) is thus also perfectly suitable for evaluating the DM comoving number density $Y_{X}$ in FI and SW scenarios for which the loss terms for the dark matter particle $X$ are neglected. [^12] For further discussion on kinetic decoupling, see e.g. [279, 280, 34]. In addition, the semi-classical Boltzmann equation (35) neglects various quantum and thermal effects, such as quantum coherence, quantum statistics, screening, multiple scatterings, and particle interactions with coherent condensates. A systematic framework for deriving quantum kinetic equations that incorporate these effects in relic density computations is provided by the Schwinger-Keldysh formalism (see, e.g., [281, 282, 283, 284, 285, 286, 287], as well as [288, 289, 290, 291] for related computations in equilibrium and [286] for a lattice-based approach). In particular for freeze-in scenarios quantum and thermal effects can alter the relic density significantly. This was demonstrated in [289] for a Majorana DM model and in [284] for a scalar DM model, respectively. The latter work, which employs non-equilibrium quantum field theory from first principles, quantifies the expected corrections to various approaches commonly used in the literature (freeze-in with only decays, including scattering with and without thermal masses, hard thermal loop approximation, etc.) and provides recommendations for accurate results in phenomenological studies. For freeze-out scenarios and in particular for co-annihilating configurations, significant impact is expected from the Sommerfeld effect and bound state formation, as discussed in section IV.2.5. However, only small corrections are expected due to thermal effects, as detailed in the context of dark matter annihilation in, e.g., [292, 281] and Sommerfeld and bound state formation in, e.g., [293, 288, 294, 295, 296, 297, 298].
@@
[…]
@@ IV.2.1 Canonical freeze-out
[…]
@@
Figure 38: Representative LO Feynman diagrams contributing to the DM annihilation cross section $\langle\sigma_{\text{eff}}v\rangle$. They include $t$-channel DM annihilations (top left), XYXY co-annihilations (top central and right), and mediator pair-annihilations (bottom row) with the blob accounting for channels yielding SM electroweak bosons or gluons. The final state $q$ represents any quark flavour, and may include a sum over flavours depending on the model. For leptophilic models, $q$ can be replaced by a lepton $\ell$, and for models with self-conjugate DM, additional YYYY annihilation diagrams must be added. Moreover, in cases where the S-wave contribution to $\langle\sigma_{\text{ann}}v\rangle(X\bar{X}\to q\bar{q})$ is helicity suppressed, NLO corrections become relevant (see the diagrams in figure 40).
@@ IV.2.2 Conversion-driven freeze-out
[…]
@@ IV.2.3 Freeze-in
Interactions with the thermal plasma can play a crucial role in scenarios of freeze-in dark matter production. This was first demonstrated in [289] for a Majorana DM model, and later extended to scalar DM models in [284, 306]. The latter works employed a first-principles non-equilibrium quantum field theory approach, specifically the closed-time path formalism, to quantify the corrections to standard approximations commonly used in the literature for freeze-in dark matter. These approximations include vacuum decays only, decays with thermal masses, decays and scatterings with thermal masses, and the hard thermal loop approximation. It was shown that using decays with vacuum masses alone can underestimate the relic abundance by up to 90% in scenarios with small mass splittings where scattering processes, particularly those involving multiple soft scatterings, dominate. In contrast, for large mass splittings where decay processes dominate, the relic abundance is predicted correctly to within $\pm\mathcal{O}(10\%)$, the sign depending on the gauge coupling. For large gauge couplings, the relic abundance is typically underestimated by about 10%, whereas for small gauge couplings it is overestimated by a similar margin. For further details, we refer the reader to [284, 306]. Scattering processes such as $VY\to Xf$ may contribute slightly to the relic abundance, up to $\sim 25\%$, when $V$ is a gluon and $f$ is a coloured SM fermion. A more precise treatment of such thermal effects, particularly for small mass splittings, may nevertheless be necessary [289]. For simplicity, we focus here on mediator decay contributions during a radiation-dominated era to illustrate the main aspects of freezing-in DM production. In this case, (42) reduces to
@@ IV.2.5 Non-perturbative effects
The freeze-out of dark sector annihilations typically occurs when the annihilating particles are non-relativistic. This regime allows for significant non-perturbative effects, namely the Sommerfeld Effect (SE) and Bound State Formation (BSF), that affect the annihilation cross section through multiple exchanges of light states between the initial or final-state particles. Non-perturbative corrections in the final state are neglected here, since the annihilation products are much lighter than the dark sector particles so that they are relativistic.
[…]
Simplified $t$-channel DM models always involve a dark sector particle $Y$ carrying SM charges. Consequently, annihilation involving two SM-charged dark sector particles is subject to corrections from multiple exchanges of (massless) gauge bosons. Each exchange of a massless gauge boson contributes parametrically with a factor of $\alpha/v$ where $\alpha$ is the relevant coupling constant and $v$ is the relative velocity between the two initial-state particles, which could be sizeable if $\alpha\sim v$. During the freeze-out of DM annihilations, $v\sim 0.1$. This thus necessitates the resummation of the ladder diagrams shown in figure 39 if $\alpha\sim 0.1$, which is particularly relevant for mediator particles $Y$ charged under $SU(3)_{C}$ (as in the models considered). In the non-relativistic regime, this resummation reduces to solving the Schrödinger equation for the wave functions of the initial-state particles, incorporating a colour potential as discussed in [293, 307, 308, 309, 310]. Since gluons are massless, the potential at energy scales far above the QCD confinement scale is Coulomb-like, and can be either attractive or repulsive depending on the representation of the initial state particles. In the context of simplified $t$-channel DM models, the dark sector particles are either singlet or lie in the (anti-)fundamental representation of $SU(3)_{C}$. Dark sector annihilation therefore leads to the following initial state colour configurations,
@@
where $\langle\sigma_{\text{BSF}}v\rangle$ is the thermally averaged cross section for bound state formation under the emission of a gluon [301]. The prefactor depends on the bound state decay and ionisation rates $\Gamma_{\text{dec}}$ and $\Gamma_{\text{ion}}$, and reflects that only bound state decays into SM particles deplete the dark sector. When excited bound states are included, transitions between different states $i$ and $j$ with rates $\Gamma_{\text{trans}}^{i\rightarrow j}$ must be considered, which yields [314, 258, 261]
@@
When summing the bound-state formation contributions to the cross section over all possible bound states in the final state, unitarity violation however arises [315, 316], although a possible solution has been recently proposed in [317].
@@
The Sommerfeld effect and bound state formation have been extensively studied in $t$-channel scenarios [288, 318, 303, 301, 304, 302, 294, 295, 258]. Their impact on the interpretation of the constraints existing in the parameter space is in particular explored in [302] for freeze-out scenarios and in [258] for CDFO models.
@@ IV.2.6 Altered cosmological histories
As shown in section III.3, these findings modify the predictions for LLP signatures, demonstrating the intricate dependence of the results on the reheating potential and the main inflaton decay channel. This arises from the power-law dilution of the relic abundance,
$Y_{X}\sim\Gamma_{Y}\times\begin{cases}\left(\dfrac{T_{\text{rh}}}{M_{Y}}\right)^{4k-1}\,&\text{Bosonic Reheating,}\\ \left(\dfrac{T_{\text{rh}}}{M_{Y}}\right)^{\frac{9-k}{k-1}}\;&\text{Fermionic Reheating,}\end{cases}$ (55)
which is true if $M_{Y}\gtrsim T_{\text{RH}}$. Although dilution of the relic abundance necessitates larger mediator decay widths in freeze-in scenarios, it reduces the DM-SM interaction strength for thermal freeze-out production. Since freeze-out during reheating has not been explored in the context of simplified $t$-channel DM models, we refer instead the reader to general discussions in e.g. [277, 328, 329].
@@ IV.3 Dark matter constraints
In this section, we describe the most relevant astrophysical and cosmological constraints that could be applied to the models studied.
In sections IV.3.1 and IV.3.2, we focus on astrophysical searches probing DM interactions at the present time, where DM is cold and clustered within halos in galaxies and galaxy clusters. The typical velocities $v$ involved range from approximately $10^{-5}c$ to $10^{-2}c$, with $v\sim 10^{-3}c$ in the Milky Way, and current detection methods primarily include indirect and direct detection approaches. Indirect detection relies on the annihilation of DM into SM particles in regions with high DM density, such as the Galactic Centre and dwarf spheroidal galaxies (dSphs). The SM particles produced through DM annihilations subsequently decay, shower, and hadronise to produce a flux of stable cosmic rays ($\gamma$, $e^{+}$, $\bar{p}$, and $\nu$), which can be detected by space-based and ground-based telescopes. Direct detection, on the other hand, relies on observing elastic scattering of DM off protons or neutrons within underground detectors, where the scattering produces a measurable recoil of the associated nucleus. However, these detection methods are most effective in probing the canonical freeze-out regime, where the DM annihilation or DM-nucleon scattering rates are sufficiently large. In contrast, scenarios that involve the freeze-in, superWIMP, or conversion-driven freeze-out mechanisms predict extremely small rates that are typically not testable with these methods. Instead, such scenarios often give rise to the interesting long-lived particle signals at colliders that have been discussed in section III.3.